Required information The following information applies to the questions displayed below. XYZ Corporation has a deferred compensation plan under which it allows certain employees to defer up to 40 percent of their salary for five years. For purposes of this problem, ignore payroll taxes in your computations. (Round your intermediate calculations and final answers to the nearest whole dollar amount.) a. Assume XYZ has a marginal tax rate of 21 percent for the foreseeable future and earns an after-tax rate of return of 16 percent on its assets. Joel Johnson, XYZs VP of finance, is attempting to determine what amount of deferred compensation XYZ should be willing to pay in five years that would make XYZ indifferent between paying the current salary of $18,400 and paying the deferred compensation. What amount of deferred compensation would accomplish this objective? Deferred compensationb. Assume Julie, an XYZ employee, has the option of participating in XYZs deferred compensation plan. Julies marginal tax rate is 37 percent and she expects the rate to remain constant over the next five years. Julie is trying to decide how much deferred compensation she will need to receive from XYZ in five years to make her indifferent between receiving the current salary of $18,400 and receiving the deferred compensation payment. If Julie takes the salary, she will invest it in a taxable corporate bond paying interest at 8 percent annually (after taxes). What amount of deferred compensation would accomplish this objective? 2 Ctiveondof eferred compensation
Answer #1

Part A

Deferred compensation $38647

After taxes= 18400*(1-21%) = 14536

Equivalent amount in five years = 14536*(1.16^5) = 30531

Deferred compensation=30531/(1-21%) = 38647

Part B

Deferred compensation $27035

After taxes = 18400*(1-37%) = 11592

Equivalent amount in five years= 11592*(1.08^5) = 17032

Deferred compensation= 17032/(1-37%) = 27035